40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  OMC elog  Not logged in ELOG logo
Entry  Tue Sep 8 11:18:10 2015, Koji, Optics, Characterization, PBS Transmission measurement setup.JPGCaF2Prism.jpgCaF2Prism2.JPG
    Reply  Wed Sep 9 01:58:34 2015, Koji, Optics, Characterization, PBS Transmission measurement 
       Reply  Thu Sep 10 04:03:42 2015, Koji, Optics, Characterization, More polarizer optics measurement (Summary) eLIGO_PBS.pdfHWP.pdfTFP.pdfGlasgow_PBS.pdf
          Reply  Wed Sep 23 17:49:50 2015, Koji, Optics, Characterization, More polarizer optics measurement (Summary) Glasgow_PBS_spotsize.pdf
Message ID: 241     Entry time: Tue Sep 8 11:18:10 2015     Reply to this: 242
Author: Koji 
Type: Optics 
Category: Characterization 
Subject: PBS Transmission measurement 

Motivation: Characterize the loss of the Calcite Brewster PBS.

Setup: (Attachment 1)

- The beam polarization is rotated by an HWP
- The first PBS filters out most of the S pol
- The second PBS further filters the S and also confirms how good the polarization is.

- The resulting beam is modulated by a chopper disk. The chopping freq can be 20~1kHz.

- The 50:50 BS splits the P-pol beam into two. One beam goes to the reference PD. The other beam goes to the measurement PD.

- Compare the transfer functions between RefPD and MeasPD at the chopping frequency with and without the DUT inserted to the measurement pass.

- The PBS shift the beam significantly. The beam can't keep the alignment on the Meas PD when the crystal is removed.
  Therefore the "On" and "Off" states are swicthed by moving the PBS and the steering mirror at the same time.
  The positions and angles of the mounts are defined by the bases on the table. The bases are adjusted to have the same spot position for these states as much as possible.

Device Under Test:

Brewster polarizer https://dcc.ligo.org/LIGO-T1300346

The prisms are aligned as shown in Attachment 2

Between the prisms, a kapton sheet (2MIL thickness) is inserted to keep the thin air gap between them.

Result:

Set1: (~max power without hard saturation)
PD1(REF) 10dB Gain (4.75kV/A) 6.39V
PD2(PBS) 10dB Gain (4.75kV/A) Thru 4.77V, PBS 4.75
Chopping frequency 234Hz, FFT 1.6kHz span AVG 20 (1s*20 = 20s)

Thru 0.748307, PBS 0.745476 => 3783 +/- 5 ppm loss
Thru 0.748227, PBS 0.745552 => 3575 +/- 5 ppm
Thru 0.748461, PBS 0.745557 => 3879 +/- 5 ppm
Thru 0.748401, PBS 0.745552 => 3806 +/- 5 ppm
Thru 0.748671, PBS 0.745557 => 4159 +/- 5 ppm
=> Loss 3841 +/- 2 ppm

Set2: (half power)
PD1(REF) 10dB Gain (4.75kV/A) 3.20V
PD2(PBS) 10dB Gain (4.75kV/A) Thru 2.38V, PBS 2.37
Chopping frequency 234Hz, FFT 1.6kHz span AVG 20 (1s*20 = 20s)

Thru 0.747618, PBS 0.744704 => 3898 +/- 5 ppm loss
Thru 0.747591, PBS 0.744690 => 3880 +/- 5 ppm
Thru 0.747875, PBS 0.744685 => 4265 +/- 5 ppm
Thru 0.747524, PBS 0.744655 => 3838 +/- 5 ppm
Thru 0.747745, PBS 0.744591 => 4218 +/- 5 ppm
=> Loss 4020 +/- 2 ppm

Set3: (1/4 power)
PD1(REF) 10dB Gain (4.75kV/A) 1.34V
PD2(PBS) 10dB Gain (4.75kV/A) Thru 1.00V, PBS 0.999
Chopping frequency 234Hz, FFT 1.6kHz span AVG 20 (1s*20 = 20s)

Thru 0.745140, PBS 0.741949 => 4282 +/- 5ppm loss
Thru 0.745227, PBS 0.741938 => 4413 +/- 5ppm
Thru 0.745584, PBS 0.741983 => 4830 +/- 5ppm
Thru 0.745504, PBS 0.741933 => 4790 +/- 5ppm
Thru 0.745497, PBS 0.741920 => 4798 +/- 5ppm
Thru 0.745405, PBS 0.741895 => 4709 +/- 5ppm
=> Loss 4637 +/- 2ppm


Possible improvement:

- Further smaller power
- Use the smaller gain as much as possible
- Compare the number for the same measurmeent with the gain changed

- Use a ND Filter instead of HWP/PBS power adjustment to reduce incident S pol
- Use a double pass configuration to correct the beam shift by the PBS

To be measured

- Angular dependence
- aLIGO Thin Film Polarizer
- HWP
- Glasgow PBS

Attachment 1: setup.JPG  1.007 MB  | Hide | Hide all
setup.JPG
Attachment 2: CaF2Prism.jpg  404 kB  Uploaded Tue Sep 8 15:01:09 2015  | Hide | Hide all
CaF2Prism.jpg
Attachment 3: CaF2Prism2.JPG  731 kB  Uploaded Tue Sep 8 15:01:23 2015  | Hide | Hide all
CaF2Prism2.JPG
ELOG V3.1.3-